# Improving Your Model Ranking on Chatbot Arena by Vote Rigging

### v1

@@ Abstract
[…]
[ backgroundcolor=newblue, linecolor=blue!80!black, linewidth=0pt, roundcorner=10pt, innertopmargin=5pt, innerbottommargin=5pt, innerleftmargin=5pt, innerrightmargin=5pt ]elegantbox
@@ 3 Vote Rigging Strategies
  • For each new vote between the sampled models $m_{a}$ and $m_{b}$, the vote manipulation function $\mathcal{M}(\widetilde{a},\widetilde{b})$ takes the identities $\widetilde{a}$ and $\widetilde{b}$ predicted by $\mathcal{A}$ as inputs and returns one of four voting options: $\widetilde{a}$/$a$ wins, $\widetilde{b}$/$b$ wins, Tie, or Abstain. Note that $\mathcal{M}$ may also depend on additional information, such as historical votes or ranks, as described in our omnipresent rigging strategy.
@@ 5 On Exploring the Rigging Capability
[…]
@@ 6.1 Towards Simulating Real-world Vote Rigging
[…]
@@
[…]
@@ 6.2 Ablation Studies on Omnipresent Rigging
Rigging the length-control leaderboard. In addition to the original leaderboard, the Chatbot Arean offers the length-control version, which explicitly disentangles the effect of response length in rating calculation. Here we aim to investigate whether our vote rigging is still effective when conducted on the length-control leaderboard. As illustrated in Table 5, our omnipresent rigging still maintains an effective rigging effect achieving an average of 4-rank improvement. Furthermore, we observe an intriguing phenomenon in which Vicuna-7B’s ranking is significantly boosted compared to rigging the original leaderboard. While our initial strategies are not specifically designed to achieve this, our findings highlight another vulnerability within the length-control mechanism, where adversaries could optimize prompts to reduce length discrepancies between responses, thereby increasing the importance of rigged votes in updating the length-control leaderboard.
@@ 7 Defense against Vote Rigging
Vote filtering with pairwise win rates. Since practical adversaries may have multiple accounts to rig which reduces the effectiveness of anomaly user detections. To address this issue, we propose a simple vote filtering designed to remove anomalous votes that deviate from the historical win rate: For each collected vote, if it satisfies $\mathcal{W}(\mathbf{r}^{\textrm{BT}}_{\mathbb{V}_{H}}[a],\mathbf{r}^{\textrm{BT}}_{\mathbb{V}_{H}}[b])>\tau$ with $m_{b}$ wins or $\mathcal{W}(\mathbf{r}^{\textrm{BT}}_{\mathbb{V}_{H}}[b],\mathbf{r}^{\textrm{BT}}_{\mathbb{V}_{H}}[a])>\tau$ with $m_{a}$ wins, we then discard it for leaderboard updating. The $\tau$ controls the filtering proportion and the basic intuition here is to reduce unlikely voting results calibrated by pairwise win rates. We provide a detailed implementation in Appendix E. As shown in Table 8, while vote filtering reduces the overall ranking improvement, it still suffers from completely eliminating the rigging effect, where our omnipresent strategies still achieve over 6-rank improvement even with $\tau=0.7$. In conclusion, our findings demonstrate the difficulty of thoroughly defending against vote rigging, implying that more effective defenses should be developed to improve the Chtbot Arena’s integrity.
Table 7: The Top-$1$ and Top-$5$ accuracy against various scales of the training corpus.
@@ 8 Related Work
LLM evaluation. Developing LLMs benchmarking is a crucial task for measuring their intrinsic capabilities. Conventional benchmarks like GLUE (Wang et al., 2018), HumanEval (Chen et al., 2021), MMLU (Hendrycks et al., 2020), and GSM-8K (Cobbe et al., 2021) assess LLMs in a static manner, where they typically rely on predefined test cases. Although convenient, these benchmarks are difficult to comprehensively capture the open-ended generation capabilities (Liang et al., 2023; Peng et al., 2022) of emerging advanced models, and are typically associated with concerns such as dataset contamination (Yang et al., 2023; Sainz et al., 2023) and Out-Of-Distribution robustness (Yuan et al., 2023). To address these challenges, recent progresses (Zheng et al., 2023a; Li et al., 2023; Dubois et al., 2024) employ LLM-as-a-Judge where a strong language model such as GPT-4 (Achiam et al., 2023) serves as a referee for model assessment. While reducing the need for human annotation, these automatic evaluators might suffer from spurious features, such as verbosity and position bias (Dubois et al., 2024; Chen et al., 2024). Unlike traditional benchmarks, Chatbot Arena (Chiang et al., 2024) devises an online platform that allows site users to vote between a pair of anonymous models based on preferred responses. By leveraging crowdsourced voting, the leaderboard aggregates high-diversity human-annotated votes, which features Chatbot Arena the most popular and widely recognized LLM benchmark.
Table 8: The mitigation results of vote filtering strategy.
@@ 9 Conclusion
Impact Statements
Due to Chatbot Arena’s widespread popularity in LLM evaluation, it is possible that practical adversaries could exploit our rigging strategies to improve their own target models’ ranking for substantial promotional benefits. These malicious behaviors would put other normal model developers’ interests at risk, and even worse, undermine the reliability and trustworthiness of Chatbot Arena. On the other hand, while we have discussed several defense methods against vote rigging, our initial attempts at rigging defense highlight the difficulties in completely eliminating the manipulation effect. As a result, we encourage the community to focus on developing more robust defense mechanisms to mitigate the rigging vulnerabilities of Chatbot Arena as well as strengthening the integrity of emerging voting-based evaluation systems such as Copilot Areana and WebDev Arena.
Appendix A Ablation Studies of Different Rigging Objectives for Omni Rigging
@@ A.2 Does Improving Relative Rating Increase Better than Improving Absolute Rating Increase for Omni-BT
In this section, we illustrate why we choose $\mathcal{R}^{\textrm{BT}}(\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}})=\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}}[t]-\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}}[\widehat{t}]$ that measures the relative rating increase between $m_{t}$ and $m_{\widehat{t}}$ as our rigging objective. For comparison, we implement a straightforward objective $\mathcal{R}^{\textrm{BT}}(\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}})=\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}}[t]$ that directly maximizes the absolute rating increase. We reconduct experiments under the setting in Section 4 and present comparison results of their average ranking increase across all manipulated battles in Figure 5. It is observed that by maximizing the relative rating increase, we achieve a more stable and efficient ranking promotion. In practice, the adversary may explore more effective rigging objectives, which is worth discussing in future studies.
@@ A.3 What if Maximizing the Win Rate of One Model for Omni-On
We experiment against six target models used in Section 4 and observe that applying $\mathcal{R}^{\textrm{On-Max}}$ results in an average $7$ ranking increase. However, switching to $\mathcal{R}^{\textrm{On-Min}}$ yields much inferior and unstable rigging results, even lowering the average ranking by 4. These findings validate the effectiveness of using the average win rate as our objective for a more stable rigging performance.
@@ B.1 Mechanisms of De-Anonymizing Functions $\mathcal{A}_{\textrm{omni}}(\cdot)$
[…]
Discussions on strategies to identify LLM through model responses. Identifying the source of LLM through model responses has been widely researched. Active identification methods such as LLM watermarking (Kirchenbauer et al., 2023; Yoo et al., 2023; Fernandez et al., 2023; Christ et al., 2024; Kirchenbauer et al., 2024) and LLM backdoor (Shu et al., 2023; Li et al., 2024d; Hubinger et al., 2024; Yan et al., 2024; Xu et al., 2024; Rando & Tramèr, 2023) embed traceable information into LLM responses, facilitating further detection through predefined statistical metrics. On the other hand, passive strategies (Dou et al., 2022; Guo et al., 2023; Chen et al., 2023a; Ghosal et al., 2023; Chen et al., 2023b; Verma et al., 2024) analyzes hidden text style without altering the generation process. For instance, Guo et al. (2023) fine-tuned a RoBERTa-based model (Liu et al., 2019) to distinguish between output preference for LLM-generated responses and human-written documents.
@@ B.2 Details of Model Selection in the Case Study
Here we elaborate on the models utilized in our case study in Section 6. For classifier training, we adopt $25$ models in total including: Llama-3-8B-Instruct (Dubey et al., 2024), Llama-2-7B-Chat, Llama-2-13B-Chat (Touvron et al., 2023), Mistral-7B-Instruct, Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Command-r (Cohere, 2024), Gemma-2B-it (Gemma et al., 2024a), Gemma-2-9B-it, Gemma-2-27B-it (Gemma et al., 2024b), Phi-3-small-8k-Instruct, Yi-34B-Chat, Yi-1.5-34B-Chat (Young et al., 2024), Qwen1.5-7B-Chat, Qwen1.5-14B-Chat (Bai et al., 2023), Starling-LM-7B-alpha, Starling-LM-7B-beta (Zhu et al., 2024), Chatglm3-6B (GLM et al., 2024), Zephyr-7B-alpha, Zephyr-7B-beta (Tunstall et al., 2023), Openchat-3.5 (Wang et al., 2023), Vicuna-7B-v1.3 (Chiang et al., 2023b), Mpt-7B-Chat (Mosaic, 2023), Wizardlm-13B (Xu et al., 2023), Solar-10.7B-instruct-v1.0 (Kim et al., 2023), and GPT-4o-mini-2024-07-18 (Chen et al., 2023a). Additionally, we involve five extra models to simulate newly added models that fall outside our initial classification range including GPT-3.5-turbo-0613 (Brown et al., 2020), Vicuna-13B-v1.3 (Chiang et al., 2023b), Phi-3-medium-4k-Instruct (Abdin et al., 2024), Dolphin-2.2.1-Mistral-7B (Hartford, 2023), and Openhermes-2.5-Mistral-7B (Teknium, 2023). In Figure 7, we demonstrate the classification performance against responses generated with unseen prompts from the HC3 and Quora datasets respectively. We evaluate each model using 1,000 responses generated from unseen prompts.
@@ C.2 Rigging Results with Additional Target Models
We provide additional comparisons of rigging performance between the target-only and omnipresent rigging strategies against 22 diverse models $m_{t}$ used in Huang et al. (2025). Our experimental results in Table 10 demonstrate the effectiveness of our omnipresent strategies, which achieve over double the ranking promotion compared to the target-only rigging strategies.
@@ Appendix D Example of Different Model Responses
We provide examples of prompts from the HC3 dataset along with corresponding model responses generated by querying model APIs on the Chatbot Arena (side by side). We show the model responses of Llama-3-8B-Instruct in Figure 8 GPT-4o-mini-2024-07-18 in Figure 9 and Gemma-2-9B-it in Figure 10 respectively.
@@ Appendix E Details of Vote Filtering
We provide a detailed description of our vote filtering strategy in Algorithm 1.

### v2

@@ 3 Vote Rigging Strategies
[…]
@@ 5 On Exploring the Rigging Capability
However, practical vote rigging is typically conducted with limited adversary’s accessibility, potentially reducing the manipulation effectiveness. Here, we conduct a series of stress tests to explore whether our strategies remain effective against these more demanding rigging scenarios. Specifically, in Section 5.1 we conduct rigging under the threat model $\{\mathbb{V}_{H},\texttt{Anonymous},\mathcal{P}_{\mathbb{V}},\emptyset\}$ to explore the impact of inaccurate de-anonymization with predicted probability $P(\widetilde{k}=k)<1$, then in Section 5.2 we conduct rigging under the threat model $\{\mathbb{V}_{H},\texttt{Real-name},\texttt{Unknown},\emptyset\}$ to simulate the influence of Unknown sampling distribution, and finally in Section 5.3, we conduct rigging under the threat model $\{\mathbb{V}_{H},\texttt{Real-name},\mathcal{P}_{\mathbb{V}},\mathbb{V}_{O}\}$ to incorporate the influence of concurrent user voting.
@@ 6.1 Towards Simulating Real-world Vote Rigging
Table 4: Rigging with prompts from the Quora (Q) and HC3 (H) datasets. We denote $\textrm{Target-Only}^{*}$ as the idealized T-Abstain.
@@
Setup for omnipresent rigging. We fine-tune RoBERTa-based classifiers (Liu et al., 2019) to classify all 25 models with two datasets respectively, including the HC3 and Quora datasets. We generate the training corpus by querying each model with 4,000 training prompts. The fine-tuning process includes 20 epochs with a batch size of 64 which takes a few hours on $\textrm{2}\times$ NVIDIA A100 GPUs. When simulating the vote rigging, we reuse the training prompts to query model pairs within each sampled battle. We defer more details of the training corpus to Appendix B.3.
@@ 6.2 Ablation Studies on Omnipresent Rigging
Rigging the length-control leaderboard. In addition to the original leaderboard, the Chatbot Arena offers the length-control version, which explicitly disentangles the effect of response length in rating calculation. Here we aim to investigate whether our vote rigging is still effective when conducted on the length-control leaderboard. As illustrated in Table 5, our omnipresent rigging still maintains an effective rigging effect achieving an average of 4-rank improvement. Furthermore, we observe an intriguing phenomenon in which Vicuna-7B’s ranking is significantly boosted compared to rigging the original leaderboard. While our initial strategies are not specifically designed to achieve this, our findings highlight another vulnerability within the length-control mechanism, where adversaries could optimize prompts to reduce length discrepancies between responses, thereby increasing the importance of rigged votes in updating the length-control leaderboard.
@@ 7 Defense against Vote Rigging
Vote filtering with pairwise win rates. Since practical adversaries may have multiple accounts to rig which reduces the effectiveness of anomaly user detections. To address this issue, we propose a simple vote filtering designed to remove anomalous votes that deviate from the historical win rate: For each collected vote, if it satisfies $\mathcal{W}(\mathbf{r}^{\textrm{BT}}_{\mathbb{V}_{H}}[a],\mathbf{r}^{\textrm{BT}}_{\mathbb{V}_{H}}[b])>\tau$ with $m_{b}$ wins or $\mathcal{W}(\mathbf{r}^{\textrm{BT}}_{\mathbb{V}_{H}}[b],\mathbf{r}^{\textrm{BT}}_{\mathbb{V}_{H}}[a])>\tau$ with $m_{a}$ wins, we then discard it for leaderboard updating. The $\tau$ controls the filtering proportion and the basic intuition here is to reduce unlikely voting results calibrated by pairwise win rates. We provide a detailed implementation in Appendix E. As shown in Table 8, while vote filtering reduces the overall ranking improvement, it still suffers from completely eliminating the rigging effect, where our omnipresent strategies still achieve over 6-rank improvement even with $\tau=0.7$. In conclusion, our findings demonstrate the difficulty of thoroughly defending against vote rigging, implying that more effective defenses should be developed to improve Chatbot Arena’s integrity.
Table 7: The Top-11 and Top-55 accuracy against various scales of the training corpus.
@@ 8 Related Work
LLM evaluation. Developing LLMs benchmarking is a crucial task for measuring their intrinsic capabilities. Conventional benchmarks like GLUE (Wang et al., 2018), HumanEval (Chen et al., 2021), MMLU (Hendrycks et al., 2020), and GSM-8K (Cobbe et al., 2021) assess LLMs in a static manner, where they typically rely on predefined test cases. Although convenient, these benchmarks are difficult to comprehensively capture the open-ended generation capabilities (Liang et al., 2023; Peng et al., 2022) of emerging advanced models, and are typically associated with concerns such as dataset contamination (Yang et al., 2023; Sainz et al., 2023) and Out-Of-Distribution robustness (Yuan et al., 2023). To address these challenges, recent progresses (Zheng et al., 2023a; Li et al., 2023; Dubois et al., 2024) employ LLM-as-a-Judge where a strong language model such as GPT-4 (Achiam et al., 2023) serves as a referee for model assessment. While reducing the need for human annotation, these automatic evaluators might suffer from spurious features, such as verbosity and position bias (Dubois et al., 2024; Chen et al., 2024; Liu et al., 2024). Unlike traditional benchmarks, Chatbot Arena (Chiang et al., 2024) devises an online platform that allows site users to vote between a pair of anonymous models based on preferred responses. By leveraging crowdsourced voting, the leaderboard aggregates high-diversity human-annotated votes, which features Chatbot Arena the most popular and widely recognized LLM benchmark.
Table 8: The mitigation results of the vote filtering strategy.
@@
Vulnerabilities of social voting systems. Social voting has become an effective method for gathering crowdsourced opinions, but it is often challenged due to potential vulnerabilities (Nitzan, 1985; Islam et al., 2010; Bassi, 2015), such as strategic manipulation. Gibbard (1973) analyzes the general vulnerability of preference aggregation to voting manipulation; Hartvigsen (2008) examines vulnerabilities in political elections; Peters & Veselova (2023) discusses the threat of group manipulation; and Wang et al. (2015); Feng et al. (2010) propose defense mechanisms to enhance the trustworthiness of voting systems. While these studies focus on various forms of strategic manipulation, our work introduces several unique challenges when manipulating the Chatbot Arena. Specifically, unlike previous manipulation settings (Bassi, 2015), where voting candidates are visible to voters, the sampled models in our case are initially anonymous before any rigging occurs. To address this challenge, we design and implement several effective model identification strategies, denoted as $\mathcal{A}$. Additionally, to improve the rigging efficiency in practice, we design two general-purpose rigging objectives ($\mathcal{R^{\textrm{BT}}}$ and $\mathcal{R^{\textrm{On}}}$), which significantly outperform baselines and the two most relevant prior works in terms of efficiency.
LLM identification through model responses. Identifying the source of LLM through model responses has been widely researched. Active identification methods such as LLM watermarking (Kirchenbauer et al., 2023; Yoo et al., 2023; Fernandez et al., 2023; Christ et al., 2024; Kirchenbauer et al., 2024) and LLM backdoor (Shu et al., 2023; Li et al., 2024d; Hubinger et al., 2024; Yan et al., 2024; Xu et al., 2024; Rando & Tramèr, 2023) embed traceable information into LLM responses, facilitating further detection through predefined statistical metrics. On the other hand, passive strategies (Dou et al., 2022; Guo et al., 2023; Chen et al., 2023a; Ghosal et al., 2023; Chen et al., 2023b; Verma et al., 2024) analyze hidden text style without altering the generation process. For instance, Guo et al. (2023) fine-tuned a RoBERTa-based model (Liu et al., 2019) to distinguish between output preference for LLM-generated responses and human-written documents.
@@ 9 Conclusion
Impact Statement
Due to Chatbot Arena’s widespread popularity in LLM evaluation, it is possible that practical adversaries could exploit our rigging strategies to improve their own target models’ ranking for substantial promotional benefits. These malicious behaviors would put other normal model developers’ interests at risk, and even worse, undermine the reliability and trustworthiness of Chatbot Arena. On the other hand, while we have discussed several defense methods against vote rigging, our initial attempts at rigging defense highlight the difficulties in completely eliminating the manipulation effect. As a result, we encourage the community to focus on developing more robust defense mechanisms to mitigate the rigging vulnerabilities of Chatbot Arena as well as strengthening the integrity of emerging voting-based evaluation systems such as Copilot Arena and WebDev Arena.
Appendix A Ablation Studies of Different Rigging Objectives for Omnipresent Rigging
@@ A.2 Does Improving Relative Rating Increase Better than Improving Absolute Rating Increase for Omni-BT
In this section, we illustrate why we choose $\mathcal{R}^{\textrm{BT}}(\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}})=\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}}[t]-\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}}[\widehat{t}$] that measures the relative rating increase between $m_{t}$ and $m_{\widehat{t}}$, i.e., the model that ranks one position ahead of $m_{t}$ as our rigging objective. For comparison, we implement a straightforward objective $\mathcal{R}^{\textrm{BT}}(\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}})=\mathbf{r}_{\mathbb{V}^{\prime}}^{\textrm{BT}}[t]$ that directly maximizes the absolute rating increase. We reconduct experiments under the setting in Section 4 and present comparison results of their average ranking increase across all manipulated votes in Figure 5. It is observed that by maximizing the relative rating increase, we achieve a more stable and efficient ranking promotion. In practice, the adversary may explore more effective rigging objectives, which is worth discussing in future studies.
@@ A.3 What if Maximizing the Win Rate of One Model for Omni-On
We experiment against six models used in Section 4 and observe that applying $\mathcal{R}^{\textrm{On-Max}}$ results in an average of 7 ranking increase. However, switching to $\mathcal{R}^{\textrm{On-Min}}$ yields much inferior and unstable rigging results, even lowering the average ranking by 4. These findings validate the effectiveness of using the average win rate as our objective for a more stable rigging performance.
@@ B.1 Mechanisms of De-Anonymizing Functions $\mathcal{A}_{\textrm{omni}}(\cdot)$
where $f_{\theta}(\cdot)[n]$ indicates its $n\textrm{-th}$ logit. Then the omnipresent de-anonymizing mechanism can be formulated as $\mathcal{A}_{\textrm{omni}}(s,m_{k}(s))=\arg\max_{n}f_{\theta^{*}}(m_{k}(s))[n]$.
@@ B.2 Details of Model Selection in the Case Study
Here, we elaborate on the models utilized in our case study in Section 6. For classifier training, we use a total of 25 models, including Llama-3-8B-Instruct (Dubey et al., 2024), Llama-2-7B-Chat, Llama-2-13B-Chat (Touvron et al., 2023), Mistral-7B-Instruct, Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Command-r (Cohere, 2024), Gemma-2B-it (Gemma et al., 2024a), Gemma-2-9B-it, Gemma-2-27B-it (Gemma et al., 2024b), Phi-3-small-8k-Instruct, Yi-34B-Chat, Yi-1.5-34B-Chat (Young et al., 2024), Qwen1.5-7B-Chat, Qwen1.5-14B-Chat (Bai et al., 2023), Starling-LM-7B-alpha, Starling-LM-7B-beta (Zhu et al., 2024), Chatglm3-6B (GLM et al., 2024), Zephyr-7B-alpha, Zephyr-7B-beta (Tunstall et al., 2023), Openchat-3.5 (Wang et al., 2023), Vicuna-7B-v1.3 (Chiang et al., 2023b), Mpt-7B-Chat (Mosaic, 2023), Wizardlm-13B (Xu et al., 2023), Solar-10.7B-instruct-v1.0 (Kim et al., 2023), and GPT-4o-mini-2024-07-18 (Chen et al., 2023a). Additionally, we involve five extra models to simulate newly added models that fall outside our initial classification range including GPT-3.5-turbo-0613 (Brown et al., 2020), Vicuna-13B-v1.3 (Chiang et al., 2023b), Phi-3-medium-4k-Instruct (Abdin et al., 2024), Dolphin-2.2.1-Mistral-7B (Hartford, 2023), and Openhermes-2.5-Mistral-7B (Teknium, 2023). In Figure 7, we demonstrate the classification performance against responses generated with unseen prompts from the HC3 and Quora datasets respectively. We evaluate each model using 1,000 responses generated from unseen prompts.
@@ C.2 Rigging Results with Additional Target Models
We provide additional comparisons of rigging performance between the target-only and omnipresent rigging strategies on target model $m_{t}$, which is set to be one of the 22 diverse models used in Huang et al. (2025). Our experimental results in Table 10 demonstrate the effectiveness of our omnipresent rigging strategies, which achieve over double the ranking promotion compared to the target-only rigging strategies.
@@ Appendix D Example of Different Model Responses
We provide examples of prompts from the HC3 dataset, as well as corresponding responses generated by querying model APIs on the Chatbot Arena (side by side). We show the model responses of Llama-3-8B-Instruct in Figure 8, GPT-4o-mini-2024-07-18 in Figure 9 and Gemma-2-9B-it in Figure 10 respectively.
@@ Appendix E Details of Vote Filtering
We provide a detailed description of our vote-filtering strategy in Algorithm 1.